OpenStudy (anonymous):

The walters average phone bill increased from $16.65 to $19.98........what is the percent of the increase????

OpenStudy (anonymous):

@ParthKohli ..............xD hi

OpenStudy (anonymous):

@apoorvk @homeisawsome @satellite73 someone help me pleasee

OpenStudy (anonymous):

Always with a percentage first decide what is the 'original' amount that represents 100%. 16.65 is the original amount here. First what is 1% =16.65/100 Then how many 1%'s go into 19.98 so you know what 19.98 is as a percentage. 16.65/100 * X = 19.98 X = 19.98/16.65 * 100 X= 120% percentage increase means "how much over 100%" = 20% increase.

OpenStudy (anonymous):

thankss